# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### 19 Dec 2023 v4

@@ 4.1 Absorption theorem
The third step, that a transitive operation witnesses absorption of left centers, is straightforward, see Proposition 8.3. A significant refinement, Corollary 7.10.2 in [60], shows that left centers are, in fact, ternary absorbing. An adjustment of the proof will also help us in proving Theorem 5.10.
@@ 4.2 Subdirect irredundant subpowers
  • the binary projections of $R_{i}$ are equal to $A^{2}$, and
  • a tuple in $R_{i}$ is determined by its values on any two coordinates.
@@ 5.4 Absorption and edges
  • $f$ witnesses all the ternary absorptions $B\trianglelefteq_{3}\mathbf{A}$;
  • any binary operation obtained from $f$ by identifying two arguments witnesses all the binary absorptions $B\trianglelefteq_{2}\mathbf{A}$.
@@ 8.1 Absorption theorem
for some $b_{1},\dots,b_{i},b_{1}^{\prime},\dots,b_{j}^{\prime}\in B$ where $i+j$ is not less than the arity of $f$. Therefore $f^{\mathbf{C}}(a+R,\dots,a+R,b_{1}+R,\dots,b_{i}+R)\subseteq a+R$ and, denoting $D=a+R$, we have $f^{\mathbf{C}}(D,\dots,D,C,\dots,C)\subseteq D$ (with $i$ occurrences of $C$ on the left). Similarly, we obtain $f^{\mathbf{C}}(C,\dots,C,D,\dots,D)\subseteq D$ with $j$ occurrences of $C$. It follows that the binary operation on $C$ obtained from $f^{\mathbf{C}}$ by identifying the first $j$ variables to $x$ and the rest to $y$ witnesses the nontrivial binary absorption $D\trianglelefteq_{2}\mathbf{C}$, a contradiction with the definition of a center. ∎
@@
and we pp-define $R^{\prime}\leq\mathbf{A}^{2n-2}$ by the formula
@@
It now remains to show that $R^{\prime}\cap B^{2n-2}=\emptyset$. Assuming the converse, there exist ${\mathbf{d}},{\mathbf{d}}^{\prime}\in R\cap(B^{n-1}\times A)$ such that $(d_{n},d^{\prime}_{n})\in S$. Let $E$ be the projection of $R$ onto the last coordinate after fixing all the other coordinates to $B$, that is, $E=\operatorname{proj}_{n}(R\cap(B^{n-1}\times A))$. Since $R$ was chosen inclusion minimal, we get $\operatorname{proj}_{n}(R)\subseteq\operatorname{Sg}_{\mathbf{A}}(B\cup\{e\})$ for every $e\in E$, otherwise we could replace ${\mathbf{a}}^{n}$ by a tuple ${\mathbf{b}}\in R\cap(B^{n-1}\times\{e\})$ and get a $B$-essential relation properly contained in $R$.
@@
By Lemma 8.8, there exists $n$ such that $nR$ is proper and $2nR$ is either $A^{2}$ or $nR$. In the latter case, $nR$ is the desired subdirect, proper, central, and transitive relation (as $nR+nR=nR$). Assume now that $nR=A^{2}$ and replace $R$ by $nR$. The relation $R$ is now subdirect, proper, and central, it satisfies $R+R=A^{2}$, and its left center is the same as the left center of the original relation (since the original relation was chosen to have the largest possible left center).
@@ 8.2 Subdirect irredundant subpowers
Without loss of generality assume $I=\{1,\dotsc,n-1\}$. Next, we claim that $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ is full. Indeed, otherwise this projection together with $\{1,2,\dots,n-2\}$ is also a counterexample, contradicting the minimality of $R$.
@@
We consider the subalgebra $\mathbf{F}$ of $\mathbf{A}^{A^{3}}$ formed by all the ternary term operations of $\mathbf{A}$, i.e., $F=\operatorname{Clo}_{3}(\mathbf{A})$. We set $P=\{(a,b):\operatorname{Sg}_{\mathbf{A}}(a,b)=\mathbf{A}\}$ and define an equivalence relation $\sim$ on $P$ by $(a,b)\sim(a^{\prime},b^{\prime})$ if $t(a,a,b)$ determines $t(a^{\prime},a^{\prime},b^{\prime})$ and vice versa; in other words, if $\{(t(a,a,b),t(a^{\prime},a^{\prime},b^{\prime})):t\in F\}$ is the graph of a bijection (from $A$ to $A$). Note that $\sim$ is indeed an equivalence relation on $A$ and that $(a,b)\sim(a^{\prime},b^{\prime})$ if, and only if, $t(a,b,a)$ and $t(a^{\prime},b^{\prime},a^{\prime})$ determine each other (since coordinates of operations in $F$ can be permuted) and this happens if, and only if, $t(b,a,a)$ and $t(b^{\prime},a^{\prime},a^{\prime})$ determine each other. Let $P^{\prime}$ be a subset of $P$ that contains exactly one representative from each $\sim$-class, let $I=\{(a,a,b),(a,b,a),(b,a,a):(a,b)\in P^{\prime}\}$, and finally let $\mathbf{R}\leq\mathbf{A}^{I}$ be the projection of $\mathbf{F}$ onto the set of coordinates $I$.
@@ 9.3 Edges
In (a) there is a binary term operation acting like the semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$. By Proposition 5.3 together with Proposition 5.4, the set $\{a/\theta,b/\theta\}$ is a subuniverse of $\mathbf{E}/\theta$ and thus is equal to it. By the classification of Post [55] and minimality of $\mathbf{E}/\theta$ we conclude (a). Case (b) is identical, except the operation is a ternary majority. In (c), Theorem 2.4 implies that we have the Mal’cev operation $x-y+z$ for some abelian group $\mathbf{G}$, so $\mathbf{E}/\theta$ is term equivalent to the affine Mal’cev algebra of $\mathbf{G}$. By the structure of finite abelian groups, $\mathbf{G}$ is isomorphic to a product of cyclic groups of prime power order. Since $E/\theta=\operatorname{Sg}_{\mathbf{E}/\theta}(a/\theta,b/\theta)=a/\theta+\{t\cdot b/\theta:t\in\mathbb{Z}\}$ (where $+$ and $\cdot$ are computed in $\mathbf{G}$), the set $\{t\cdot b/\theta:t\in\mathbb{Z}\}$ covers $G$, therefore the orders of any two distinct factors in the decomposition of $\mathbf{G}$ are powers of two distinct primes. We conclude that $\mathbf{G}$ is isomorphic to $\mathbb{Z}/m$ for some $m$.
@@
We are left with the case of $B\trianglelefteq_{3}\mathbf{A}$ and $R\cap D^{2}=\emptyset$. Note that $D\trianglelefteq_{3}\mathbf{C}$ and, by Theorem 5.10, $R\subseteq B(x)\vee B(y)$. Therefore $D+R=C\setminus D$ is a ternary absorbing subuniverse of $\mathbf{C}$. By item (2) in Proposition 5.12, we conclude that the partition of $C$ into $D$ and $C\setminus D$ defines a congruence of $\mathbf{C}$ and the quotient modulo this congruence is term equivalent to a two-element majority algebra. This finishes the proof. ∎
@@ 9.4 Absorption and edges
Before proving Theorem 5.19 we give a useful criterion for s-closed subsets to be binary absorbing.
@@
If $\mathbf{A}$ has a nontrivial affine quotient, then since $B$ is stable under abelian edges, $B$ must intersect the congruence class which contains $a$, so we can apply the inductive assumption to find an s-edge from $a$ to $B$ which is contained in that congruence class. Otherwise, Lemma 9.11 implies that there is some nontrivial ternary absorbing subalgebra $\mathbf{C}$ of $\mathbf{A}$.
@@
Now we have $C\cap B\neq\emptyset$ and we can apply the induction hypothesis to see that $C\cap B$ 2-absorbs $\mathbf{C}$. This implies that $C\cap B$ is a ternary absorbing subalgebra of $\mathbf{A}$ by the transitivity of 3-absorption (item (3) in Proposition 5.12). Thus we may assume without loss of generality that $C\subseteq B$.
@@
We claim that $C$ 2-absorbs the subalgebra $\mathbf{D}$ of $\mathbf{A}$ with universe $a+S$ (which contains $a$ since $S$ is reflexive). By Proposition 9.10. we just need to check that for any $d\in D\setminus C$ and any $c\in C$, $\operatorname{Sg}(d,c)$ has a proper binary absorbing subuniverse. To see this, note that the binary relation $T=S\cap(A\times\operatorname{Sg}(d,c))$ is a subdirect subuniverse of $\mathbf{A}\times\operatorname{Sg}(d,c)$ (since $c$ is in the right center of $S$) and $a$ is contained in the left center of $T$ (since $d,c\in a+S$). Since $d\not\in C$, we see that $T$ is a proper subset of $A\times\operatorname{Sg}(d,c)$, so the left center of $T$ is a proper subalgebra of $\mathbf{A}$. Thus if $\operatorname{Sg}(d,c)$ has no proper binary absorbing subuniverse, then the left center of $T$ is a center of $\mathbf{A}$, and $a$ is contained in a proper ternary absorbing subuniverse of $\mathbf{A}$ by Proposition 4.3, a contradiction. The same argument shows that $C$ 2-absorbs a subalgebra containing $b$. Application of Lemma 9.15 then gives us a contradiction, so we have proved that every reflexive subdirect subuniverse of $\mathbf{A}^{2}$ is the equality or the full relation.
@@
Finally, consider the subdirect symmetric ternary relation $R=\operatorname{Sg}_{\mathbf{A}^{3}}\{(a,a,b),(a,b,a),(b,a,a)\}.$ Since the projection $\operatorname{proj}_{12}(R)$ of $R$ onto the first two coordinates contains $(a,a),(a,b)$ and $\operatorname{Sg}(a,b)=A$, we see that $\operatorname{proj}_{12}(R)$ has a left center that contains $a$. The center cannot be proper since otherwise we get a nontrivial 2-absorbing subuniverse or the center of $\operatorname{proj}_{12}(R)$ is a Taylor center of $\mathbf{A}$ and we apply Proposition 4.3 to get a contradiction. Therefore $\operatorname{proj}_{12}(R)=A^{2}$. If $(a,a,a)\in R$, then the same argument as in Lemma 9.11 shows that $(a,b)$ is a majority edge, which gives us a proper ternary subuniverse containing $a$ (and also one containing $b$), a contradiction. The binary relation $T(x,y)\equiv R(a,x,y)$ is thus subdirect (as the projection of $R$ onto any two coordinates is full) but not full (as $(a,a)\not\in T$). Observe that $T+S$ is reflexive but cannot be full (as otherwise $T=(T+S)+S$ is full as well), therefore it is the equality relation and hence $T$ is a graph of a bijection. It follows that the strongly projective subuniverse $B$ from Lemma 9.12 is neither empty (as $\mathbf{A}$ is not abelian) nor equal to $A$ (as $a\not\in B$). Theorem 5.7 shows that $B$ 2-absorbs $\mathbf{A}$, which is a contradiction. ∎
@@
By Theorem 9.16, we see that either $\operatorname{Sg}_{\mathbf{A}}(a,b)$ has a nontrivial abelian quotient and then $\operatorname{Sg}_{\mathbf{A}}\{a,b\}/\theta$ can be chosen of prime order by taking a maximal $\theta$ (by Theorem 5.13), or one of $a,b$ is contained in a proper ternary absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$.
@@
In both cases, we see that there is some ternary operation $t\in\operatorname{Clo}(\mathbf{G})$ which witnesses the ternary absorption $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$, and we may suppose without loss of generality that $t$ acts like the first projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$. Then we have

### 21 Feb 2024 v5

@@ 4.1 Absorption theorem
The third step, that a transitive operation witnesses absorption of left centers, is straightforward, see Proposition 8.3. A significant refinement, Corollary 7.10.2 in [60], shows that left centers are, in fact, 3-absorbing. An adjustment of the proof will also help us in proving Theorem 5.10.
@@ 4.2 Subdirect irredundant subpowers
  • the binary projections of $R$ are equal to $A^{2}$, and
  • every tuple in $R$ is determined by its values on any two coordinates.
@@ 5.4 Absorption and edges
  • $f$ witnesses all the 3-absorptions $B\trianglelefteq_{3}\mathbf{A}$;
  • any binary operation obtained from $f$ by identifying two arguments witnesses all the 2-absorptions $B\trianglelefteq_{2}\mathbf{A}$.
@@ 8.1 Absorption theorem
for some $b_{1},\dots,b_{i},b_{1}^{\prime},\dots,b_{j}^{\prime}\in B$ where $i+j$ is not less than the arity of $f$. Therefore $f^{\mathbf{C}}(a+R,\dots,a+R,b_{1}+R,\dots,b_{i}+R)\subseteq a+R$ and, denoting $D=a+R$, we have $f^{\mathbf{C}}(D,\dots,D,C,\dots,C)\subseteq D$ (with $i$ occurrences of $C$ on the left). Similarly, we obtain $f^{\mathbf{C}}(C,\dots,C,D,\dots,D)\subseteq D$ with $j$ occurrences of $C$. It follows that the binary operation on $C$ obtained from $f^{\mathbf{C}}$ by identifying the first $j$ variables to $x$ and the rest to $y$ witnesses the nontrivial 2-absorption $D\trianglelefteq_{2}\mathbf{C}$, a contradiction with the definition of a center. ∎
@@
Note that $a^{n}_{n}$ is not in $B$. We will now work towards showing that $(a^{n}_{n},a^{n}_{n})$ is in $S$, which will contradict the assumption on $B$.
We pp-define $R^{\prime}\leq\mathbf{A}^{2n-2}$ by the formula
@@
Since $R^{\prime}\cap B^{2n-2}\neq\emptyset$, there exist ${\mathbf{d}},{\mathbf{d}}^{\prime}\in R\cap(B^{n-1}\times A)$ such that $(d_{n},d^{\prime}_{n})\in S$. Let $E$ be the projection of $R$ onto the last coordinate after fixing all the other coordinates to $B$, that is, $E=\operatorname{proj}_{n}(R\cap(B^{n-1}\times A))$. Since $R$ was chosen inclusion minimal, we get $\operatorname{proj}_{n}(R)\subseteq\operatorname{Sg}_{\mathbf{A}}(B\cup\{e\})$ for every $e\in E$, otherwise we could replace ${\mathbf{a}}^{n}$ by a tuple ${\mathbf{b}}\in R\cap(B^{n-1}\times\{e\})$ and get a $B$-essential relation properly contained in $R$.
@@
By Lemma 8.8, there exists $n$ such that $nR$ is proper and $2nR$ is either $A^{2}$ or $nR$. In the latter case, $nR$ is the desired subdirect, proper, central, and transitive relation (as $nR+nR=nR$). Assume now that $2nR=A^{2}$ and define $S=nR$. The relation $S$ is subdirect, proper, and central, it satisfies $S+S=A^{2}$, and its left center is the same as the left center of $R$ (since $R$ was chosen to have the largest possible left center).
@@ 8.2 Subdirect irredundant subpowers
Without loss of generality assume $I=\{1,\dotsc,n-1\}$. Next, we claim that $\operatorname{proj}_{1,\dotsc,n-2,n}(R)$ is full. Indeed, otherwise this projection together with $\{1,2,\dots,n-2\}$ is also a counterexample, contradicting the minimality of $n$.
@@
We consider the subalgebra $\mathbf{F}$ of $\mathbf{A}^{A^{3}}$ formed by all the ternary term operations of $\mathbf{A}$, i.e., $F=\operatorname{Clo}_{3}(\mathbf{A})$. We set $P=\{(a,b):\operatorname{Sg}_{\mathbf{A}}(a,b)=\mathbf{A}\}$ and define an equivalence relation $\sim$ on $P$ by $(a,b)\sim(a^{\prime},b^{\prime})$ if $t(a,a,b)$ determines $t(a^{\prime},a^{\prime},b^{\prime})$ and vice versa; in other words, if $\{(t(a,a,b),t(a^{\prime},a^{\prime},b^{\prime})):t\in F\}$ is the graph of a bijection (from $A$ to $A$). Note that $\sim$ is indeed an equivalence relation on $P$ and that $(a,b)\sim(a^{\prime},b^{\prime})$ if, and only if, $t(a,b,a)$ and $t(a^{\prime},b^{\prime},a^{\prime})$ determine each other (since coordinates of operations in $F$ can be permuted) and this happens if, and only if, $t(b,a,a)$ and $t(b^{\prime},a^{\prime},a^{\prime})$ determine each other. Let $P^{\prime}$ be a subset of $P$ that contains exactly one representative from each $\sim$-class, let $I=\bigcup_{(a,b)\in P^{\prime}}\{(a,a,b),(a,b,a),(b,a,a)\}$, and finally let $\mathbf{R}\leq\mathbf{A}^{I}$ be the projection of $\mathbf{F}$ onto the set of coordinates $I$.
@@ 9.3 Edges
In (a) there is a binary term operation acting like the semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$. By Proposition 5.3 together with Proposition 5.4, the set $\{a/\theta,b/\theta\}$ is a subuniverse of $\mathbf{E}/\theta$ and thus is equal to it. By the classification of Post [55] and minimality of $\mathbf{E}/\theta$ we conclude (a). Case (b) is identical, except the operation is a ternary majority. In (c), Theorem 2.4 implies that we have the Mal’cev operation $x-y+z$ for some abelian group $\mathbf{G}$, so $\mathbf{E}/\theta$ is term equivalent to the affine Mal’cev algebra of $\mathbf{G}$. Since $E/\theta=\operatorname{Sg}_{\mathbf{E}/\theta}(a/\theta,b/\theta)=a/\theta+\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ (where $+$ and $\cdot$ are computed in $\mathbf{G}$), the set $\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ covers $G$, therefore the group homomorphism $\mathbb{Z}\to\mathbf{G}$ defined by $t\mapsto t(b/\theta-a/\theta)$ is surjective. We conclude using the first isomorphism theorem that $\mathbf{G}$ is isomorphic to $\mathbb{Z}/m$ for some $m$.
@@
We are left with the case of $B\trianglelefteq_{3}\mathbf{A}$ and $R\cap D^{2}=\emptyset$. Note that $D\trianglelefteq_{3}\mathbf{C}$ and, by Theorem 5.10, $R\subseteq B(x)\vee B(y)$. Therefore $D+R=C\setminus D$ is a 3-absorbing subuniverse of $\mathbf{C}$. By item (2) in Proposition 5.12, we conclude that the partition of $C$ into $D$ and $C\setminus D$ defines a congruence of $\mathbf{C}$ and the quotient modulo this congruence is term equivalent to a two-element majority algebra. This finishes the proof. ∎
@@ 9.4 Absorption and edges
Before proving Theorem 5.19 we give a useful criterion for s-closed subsets to be 2-absorbing.
@@
If $\mathbf{A}$ has a nontrivial affine quotient, then since $B$ is stable under abelian edges, $B$ must intersect the congruence class which contains $a$, so we can apply the inductive assumption to find an s-edge from $a$ to $B$ which is contained in that congruence class. Otherwise, Lemma 9.11 implies that there is some nontrivial 3-absorbing subalgebra $\mathbf{C}$ of $\mathbf{A}$.
@@
Now we have $C\cap B\neq\emptyset$ and we can apply the induction hypothesis to see that $C\cap B$ 2-absorbs $\mathbf{C}$. This implies that $C\cap B$ is a 3-absorbing subalgebra of $\mathbf{A}$ by the transitivity of 3-absorption (item (3) in Proposition 5.12). Thus we may assume without loss of generality that $C\subseteq B$.
@@
We claim that $C$ 2-absorbs the subalgebra $\mathbf{D}$ of $\mathbf{A}$ with universe $a+S$ (which contains $a$ since $S$ is reflexive). By Proposition 9.10. we just need to check that for any $d\in D\setminus C$ and any $c\in C$, $\operatorname{Sg}(d,c)$ has a proper 2-absorbing subuniverse. To see this, note that the binary relation $T=S\cap(A\times\operatorname{Sg}(d,c))$ is a subdirect subuniverse of $\mathbf{A}\times\operatorname{Sg}(d,c)$ (since $c$ is in the right center of $S$) and $a$ is contained in the left center of $T$ (since $d,c\in a+S$). Since $d\not\in C$, we see that $T$ is a proper subset of $A\times\operatorname{Sg}(d,c)$, so the left center of $T$ is a proper subalgebra of $\mathbf{A}$. Thus if $\operatorname{Sg}(d,c)$ has no proper 2-absorbing subuniverse, then the left center of $T$ is a center of $\mathbf{A}$, and $a$ is contained in a proper 3-absorbing subuniverse of $\mathbf{A}$ by Proposition 4.3, a contradiction. The same argument shows that $C$ 2-absorbs a subalgebra containing $b$. Application of Lemma 9.15 then gives us a contradiction, so we have proved that every reflexive subdirect subuniverse of $\mathbf{A}^{2}$ is the equality or the full relation.
@@
Finally, consider the subdirect symmetric ternary relation $R=\operatorname{Sg}_{\mathbf{A}^{3}}\{(a,a,b),(a,b,a),(b,a,a)\}.$ Since the projection $\operatorname{proj}_{12}(R)$ of $R$ onto the first two coordinates contains $(a,a),(a,b)$ and $\operatorname{Sg}(a,b)=A$, we see that $\operatorname{proj}_{12}(R)$ has a left center that contains $a$. The center cannot be proper since otherwise we get a nontrivial 2-absorbing subuniverse or the center of $\operatorname{proj}_{12}(R)$ is a Taylor center of $\mathbf{A}$ and we apply Proposition 4.3 to get a contradiction. Therefore $\operatorname{proj}_{12}(R)=A^{2}$. If $(a,a,a)\in R$, then the same argument as in Lemma 9.11 shows that $(a,b)$ is a majority edge, which gives us a proper 3-absorbing subuniverse containing $a$ (and also one containing $b$), a contradiction. The binary relation $T(x,y)\equiv R(a,x,y)$ is thus subdirect (as the projection of $R$ onto any two coordinates is full) but not full (as $(a,a)\not\in T$). Observe that $T+S$ is reflexive but cannot be full (as otherwise $T=(T+S)+S$ is full as well), therefore it is the equality relation and hence $T$ is a graph of a bijection. It follows that the strongly projective subuniverse $B$ from Lemma 9.12 is neither empty (as $\mathbf{A}$ is not abelian) nor equal to $A$ (as $a\not\in B$). Theorem 5.7 shows that $B$ 2-absorbs $\mathbf{A}$, which is a contradiction. ∎
@@
By Theorem 9.16, we see that either $\operatorname{Sg}_{\mathbf{A}}(a,b)$ has a nontrivial abelian quotient and then $\operatorname{Sg}_{\mathbf{A}}\{a,b\}/\theta$ can be chosen of prime order by taking a maximal $\theta$ (by Theorem 5.13), or one of $a,b$ is contained in a proper 3-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$.
@@
In both cases, we see that there is some ternary operation $t\in\operatorname{Clo}(\mathbf{G})$ which witnesses the 3-absorption $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$, and we may suppose without loss of generality that $t$ acts like the first projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$. Then we have
@@ 11 Conclusion
Acknowledgement
The authors thank the referees for the many very good comments that helped to improve the proofs.